OpenStudy (anonymous):

HELP ME ? I WILL GIVE YOU A MEDAL & FAN The altitude of a plane as it begins to descend is shown by the linear equation a = 28,500 – 1800m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 9?

OpenStudy (anonymous):

12,300

OpenStudy (anonymous):

A=28500-(1800x9) Multiply 1800 by 9 A=28500-16200 Subtract A=12,300